OpenStudy (anonymous):

Solve -2x + 1 -7 and describe the graph of the solution.

jimthompson5910 (jim_thompson5910):

what's between the 1 and the -7?

OpenStudy (anonymous):

Shoot. < _

jimthompson5910 (jim_thompson5910):

First step is to subtract 1 from both sides. Then you divide both sides by -2 (this will flip the inequality sign) \[\Large -2x + 1 \le -7\] \[\Large -2x \le -7-1\] \[\Large -2x \le -8\] \[\Large x \ge \frac{-8}{-2}\] \[\Large x \ge 4\] How would you describe the graph of this?

OpenStudy (anonymous):

he is correct so this located on the x axis would be 4 or greater

OpenStudy (anonymous):

closed circle, shading to the left?

OpenStudy (anonymous):

correct

jimthompson5910 (jim_thompson5910):

close, but it's not to the left

jimthompson5910 (jim_thompson5910):

the \(\Large \ge\) sign tells us to shade to the right

OpenStudy (anonymous):

Thanks ♥

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

lol hes right again > you go in the directions its pointing

OpenStudy (anonymous):

I'll learn one day, lol

OpenStudy (anonymous):

sorry i couldnt have been of more assisstance this its been a year

OpenStudy (anonymous):

it's okay, :)